OpenStudy (anonymous):

There are 96 pencils packed in 8 boxes. How many boxes are needed for 180 pencils?

OpenStudy (harsha19111999):

96 pencils are packed in 8 boxes. So, number of pencils packed in 1 box = 96/8=12.

OpenStudy (harsha19111999):

Number of boxes needed for 180 pencils = 180/12 = 15 boxes

OpenStudy (anonymous):

So 15 is the final answer?

OpenStudy (harsha19111999):

Yes

OpenStudy (igreen):

96 pencils packed in 8 boxes First find out how many pencils are in 1 box. 96 / 8 = 12 Then divide 180 / 12 = 15